mk graph Manual

Edition : 2009.02.12

OpenXM.org

1 Function Manual

1.1 Outline

1.2 Notation

1.3 Functions

1.3.1 mtg.plot3d
mtg.plot3d(formula)

:: Draw a graph of formula

mtg.plot3d(formula | options)

:: Draw a graph of formula. Optional arguments are described below.

return List formula Expression or quote data. It should be a function in x, y. optinal arguments

domain List. [[xmin,xmax],[ymin,ymax]]

mesh Natural number. Division number to mesh the region. fit When it is 1, (max+min)/2 is moved to the orgin in z.

• Details have not been written. See examples.

     [1210] import("mt_graph.rr");
     [1211] mtg.test2();
     [1210] import("mt_graph.rr");
     [1211] mtg.plot3d(x^2-y^2);
     [1210] import("mt_graph.rr");
     [1211] mtg.plot3d(x^2-y^2 | domain=[[-1,1],[1,1]]);
     [1210] import("mt_graph.rr");
     [1211] def myfunc(X,Y) { if (X*Y < 0) return 0; else 1;}
     [1212] mtg.plot3d(quote(myfunc(x,y)*x*y));
1.3.2 mtp.parametric_plot3d
mtp.parametric_plot3d(formula)

:: Draw a graph of formula

mtp.parametric_plot3d(formula | options)

:: Draw a graph of formula. Optinal arguments are described below.

return List formula Expression or quote data. It should be a function in s, t. optinal arguments

domain List. [[xmin,xmax],[ymin,ymax]] mesh Natural number. Division number to mesh the region. fitting If it is set to 0, then automatic fitting to the z-direction is not done.

• Details have not been written. See examples.

[1210] import("mt_graph.rr");
[1211] mtp.test5();   /*  Klein bottle (8 figure) */
[1210] import("mt_graph.rr");
[1211] mtp.parametric_plot3d([s,t,s^2-t^2]);
[1210] import("mt_graph.rr");
[1211] def myfunc(X,Y) { if (X*Y < 0) return 0; else 1;}
[1212] mtp.parametric_plot3d([s,t,quote(myfunc(s,t)*s*t) | fitting=0);

Index 3

Index

(Index is nonexistent) (Index is nonexistent)

Short Contents

1 Function Manual. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

Table of Contents

1 Function Manual. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 Outline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Notation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.3.1 mtg.plot3d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3.2 mtp.parametric_plot3d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3